# Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search

### v1

@@ 5 Conclusion
This manuscript presented the convergence and rate-of-convergence results of two types of nonmonotine proximal gradient methods, i.e., combined with average line search and max line search, respectively. These results were established under the KL property, requiring only that the smooth part of the objective function has locally (rather than globally) Lipschitz continuous gradients. Note that the partitioning of the index set is a useful tool for the convergence of NPG. At least for NPG with average line search, the convergence theory is independent on the specific choice of partitioning. Additionally, our results in Section 4 address the question raised in [17, Final Remarks], demonstrating that the technique of proof there can indeed be extended to NPG with max line search.

### v3

@@ 5 Conclusion
This manuscript presented the convergence and rate-of-convergence results of two types of NPG methods, i.e., those combined with average line search and max line search, respectively. The results were established under the KL property, requiring only that the smooth part of the objective function has locally (rather than globally) Lipschitz continuous gradients. Note that the partitioning of the index set is a useful tool for the convergence analysis of NPG methods. At least for the NPG method with average line search, the convergence theory is independent on the specific choice of partitioning. Additionally, our results in Section 4 addressed the question raised in [18, Final Remarks], demonstrating that the technique of proof there can be extended to the NPG method with max line search.